Use the following information to answer the remaining questions: A firm sells a product in a perfectly competitive market, at a price of $6 each. The firm has a fixed cost of $12. Fill in the following table. Marginal Cost TC TR Profit Marginal Revenue Output AFC AVC 4 2 24
Answer #1

Fixed cost = $12

Price= $6

AFC= FC/Q

ATC= AFC+AVC

ATC= TC/Q

TC=FC when Q=0

TR= P(Q)

MR=Change in TR

MC= Change in TC

Profit= TR-TC

Output AFC AVC ATC TC TR MR MC Profit
0 - - - 12 0 - - -12
1 12 4 16 16 6 6 4 -10
2 6 3 9 18 12 6 2 -6
3 4 4 8 24 18 6 6 -6